OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation x - 2y = 4 using the x- and y- intercepts method ? Choose one answer. a. Plot the y-intercept (0, -2). b. Plot the intercept (4, 0). c. Count up 1 and right 2 from the y-intercept. d. Draw a line through the two points.

OpenStudy (anonymous):

This is x and y intercept method: Make x=0 to find y-intercept. x - 2y = 4 0 - 2y = 4 -2y = 4 /-2 y = -2 (0, -2) Make y=0 to find x-intercept x - 2y = 4 x - 2(0) = 4 x - 0 = 4 x = 4 (4, 0) And then cross a line through the two points. C is the answer. Why? You only do that step when you change the equation to slope-incercept and solve.

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

You're very welsome :)

OpenStudy (anonymous):

Welcome*